# The Race to Efficiency: A New Perspective on AI Scaling Laws

### v1

@@ 3.4 Deriving the Relative-Loss Equation
$L(t)$ $=\;L_{0}\Bigl{(}1+\tfrac{\Delta C(t)}{C_{0}}\Bigr{)}^{-\kappa}\;=\;L_{0}\left(1+\frac{2^{\gamma t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\right)^{-\kappa},$ (3.3)
@@ 4.1 Reduction to Classical Scaling Laws at $\gamma=0$
$R(t)\,=\,\Bigl{(}1\;+\;\tfrac{2^{\gamma\,t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\Bigr{)}^{-\kappa},\quad\text{where}\quad L(t)\;=\;L_{0}\,R(t),$
@@
$L(t)\,=\;L_{0}\,\Bigl{(}1\;+\;\tfrac{t}{1\,\text{yr}}\Bigr{)}^{-\kappa}.$
@@ 5 Implications and Case Studies
Having introduced a time- and efficiency-aware perspective on AI scaling, we illustrate its consequences through several thought experiments. First, we examine two theoretically static scenarios ($\gamma=0$), where compute is “unfolded” either in space (front-loading all GPUs simultaneously) or in time (running a fixed-size cluster for millennia). Next, we consider scenarios with positive $\gamma$, exploring how a balance between up-front GPU investment and sustained efficiency gains can shape multi-year outcomes.
@@ 5.1 How Severe Are the Diminishing Returns in AI Scaling?
$\Bigl{(}1+\tfrac{t}{1\,\text{yr}}\Bigr{)}^{-\kappa}\;=\;\Bigl{(}1+\tfrac{\Delta C}{C_{0}}\Bigr{)}^{-\kappa}\;=\;0.68,$
@@
These two extreme scenarios reveal that the main hurdle is not the $L=0.68$ target itself; even less ambitious goals fare no better under a strictly static $\gamma=0$ assumption. Consequently, $\boldsymbol{\gamma>0}$ emerges as a theoretical requirement for computational feasibility. Formally,
$L(t)\;=\;L_{0}\,\Bigl{(}1+\tfrac{2^{\gamma t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\Bigr{)}^{-\kappa},\quad\gamma>0,$
or else the small scaling exponent $\kappa$ imposes diminishing returns with prohibitive resource or time costs.
@@ 5.2 A Multi-Year Case: Baseline, Turtle, and Hare
$L_{0}\;=\;L_{\mathrm{init}}\,\bigl{(}1+\tfrac{\Delta C}{C_{0}}\bigr{)}^{-\kappa}.$
@@ 5.3 Scaling Laws as the Driving Force for Innovations
$L(t)\;=\;L_{0}\,\Bigl{(}1+\tfrac{\Delta C(t)}{C_{0}}\Bigr{)}^{-\kappa},$

### v2

@@ 3.4 Deriving the Relative-Loss Equation
$L(t)$ $=\;L_{0}\Bigl{(}1+\frac{\Delta C(t)}{C_{0}}\Bigr{)}^{-\kappa}\;=\;L_{0}\left(1+\frac{2^{\gamma t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\right)^{-\kappa},$ (3.3)
@@ 4.1 Reduction to Classical Scaling Laws at $\gamma=0$
$R(t)\,=\,\Bigl{(}1\;+\;\frac{2^{\gamma\,t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\Bigr{)}^{-\kappa},\quad\text{where}\quad L(t)\;=\;L_{0}\,R(t),$
@@
$L(t)\,=\;L_{0}\,\Bigl{(}1\;+\;\frac{t}{1\,\text{yr}}\Bigr{)}^{-\kappa}.$
@@ 5 Implications and Case Studies
Having introduced a time- and efficiency-aware perspective on AI scaling, we illustrate its consequences through several thought experiments. First, we examine two theoretically static scenarios ($\gamma=0$), where compute is “unfolded” either in space (front-loading all GPUs simultaneously) or in time (running a fixed-size cluster for millennia). Next, we consider scenarios with positive $\gamma$, exploring how a balance between up-front GPU investment and sustained efficiency gains can shape multi-year outcomes. Please see Table 2 for the numerical results of these experiments.
@@ 5.1 How Severe Are the Diminishing Returns in AI Scaling?
$\Bigl{(}1+\frac{t}{1\,\text{yr}}\Bigr{)}^{-\kappa}\;=\;\Bigl{(}1+\frac{\Delta C}{C_{0}}\Bigr{)}^{-\kappa}\;=\;0.68,$
@@
Classical scaling laws, when taken in a purely static sense ($\gamma=0$), imply that driving the loss from $L_{0}=1.0$ to $L=0.68$ might require a 3,000$\times$ increase in compute (or equivalently, 3,000 years of training on the same hardware). However, this enormous “3,000$\times$ factor” does not directly translate to realistic time and resource constraints. By introducing $\gamma>0$ (the efficiency-doubling rate), our framework extends the predictive power of classical scaling laws to more realistic settings, where hardware refreshes, architecture refinements, and data-pipeline optimizations occur continually. Mathematically, one now has
$L(t)\;=\;L_{0}\,\Bigl{(}1+\frac{2^{\gamma t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\Bigr{)}^{-\kappa},\quad\gamma>0,$
so that periodic improvements in “usable compute” offset the high cost of further reducing loss.
@@ 5.2 A Multi-Year Case: Baseline, Turtle, and Hare
$L_{0}\;=\;L_{\mathrm{init}}\,\bigl{(}1+\frac{\Delta C}{C_{0}}\bigr{)}^{-\kappa}.$
@@ 5.3 Scaling Laws as the Driving Force for Innovations
$L(t)\;=\;L_{0}\,\Bigl{(}1+\frac{\Delta C(t)}{C_{0}}\Bigr{)}^{-\kappa},$